Poly melts at approximately at 327 degrees celsius. If the heat of fusion is 5.28 kJ mol ^-1, what is the molar entropy change oc fusion?

1
Expert's answer
2020-10-09T13:46:27-0400

Temperature  T=327OC=600K
Heat, q=600×5.28kJ Entropy change, ΔS=K
600×5.8kJ​=3480KJ

(3480)/(6.023)(1023) = 5.778-21